OpenStudy (anonymous):

Identify the GCF of 6x^2y^2 − 8xy^2 + 10xy^3. A. 6x^2y^2 B. 2x^2y C. 2xy^2 D. 6xy^2

OpenStudy (anonymous):

whats the gcf of 6,10 and 8

OpenStudy (anonymous):

2

OpenStudy (anonymous):

i got 2x^2y

OpenStudy (anonymous):

correct i belive

OpenStudy (anonymous):

i'd say its 2x^y 2xy (3 xy - 4y + 5 y^2)

OpenStudy (anonymous):

but 2x^y isnt a choice...

OpenStudy (jdoe0001):

\(\bf 6x^2y^2 - 8xy^2 + 10xy^3\\ \textit{let us factor decompose each term}\\ 6x^2y^2 \implies 2\cdot 3\cdot xxyy\\ 8xy^2\implies 2\cdot 2\cdot 2 \cdot xyy\\ 10xy^3\implies 2 \cdot 5\cdot xyyy\\ \quad \\ 6x^2y^2 - 8xy^2 + 10xy^3\\\quad \\ \implies 2\cdot 3\cdot xxyy-2\cdot 2\cdot 2 \cdot xyy+2 \cdot 5\cdot xyyy\) see the GCF?

OpenStudy (anonymous):

i meant : 2x^2 y .. . :D

OpenStudy (anonymous):

thanks guys